OpenStudy (anonymous):

? What is the next step in the proof? Choose the most logical approach. Given: Transversal t passes through parallel lines r and s. Prove: ∠3≅∠6 ∠4≅∠5 Statement Reason 1. r || s given 2. ∠3≅∠7 ∠4≅∠8 For parallel lines cut by a transversal, corresponding angles are congruent. 3.

OpenStudy (anonymous):

? What is the next step in the proof? Choose the most logical approach. Statement: ∠1≅∠8 and ∠2≅∠7 Reason: Congruent Supplements Theorem Statement: m∠3+m∠4=180° and m∠7+m∠8=180° Reason: Linear Pair Theorem Statement: m∠3+m∠5=180° and m∠4+m∠6=180° Reason: definition of supplementary angles Statement: ∠7≅∠6 and ∠8≅∠5 Reason: Vertical Angles Theorem Done Given: Transversal t passes through parallel lines r and s. Prove: ∠3≅∠6 ∠4≅∠5 Statement Reason 1. r || s given 2. ∠3≅∠7 ∠4≅∠8 For parallel lines cut by a transversal, corresponding angles are congruent. 3.